# Myhill-Nerode for hypergraphs and an application to gain-graphic matroids

### v2

@@ 1. Introduction
In this article, we present a version of the Myhill-Nerode theorem for hypergraphs, which we will use to prove that certain classes of hypergraphs are not monadically definable. In particular, we will prove non-definability results when the hypergraphs are gain-graphic matroids. In the classical Myhill-Nerode theorem, $\mathcal{L}$ is a language of strings and two strings, $\mathbf{w}_{1}$ and $\mathbf{w}_{1}$, are equivalent relative to $\mathcal{L}$, if, for every possible string $\mathbf{z}$, either both of the concatenations $\mathbf{w}_{1}\mathbf{z}$ and $\mathbf{w}_{2}\mathbf{z}$ are in $\mathcal{L}$, or neither is. The language is regular if and only if this equivalence relation has finite index. To present a theorem of this type for hypergraphs, we need an operation for hypergraphs that plays the same role as concatenation in the classical theorem. Let $C$, $U$, and $V$ be finite sets where $U\cap V=\emptyset$. Let $c$ be a function from $2^{U}$ to $C$ and let $d$ be a function from $2^{V}\times C$ to $\{0,1\}$. We can glue $(U,c)$ and $(V,d)$ together to produce a hypergraph $(U,c)\boxplus(V,d)$ with the ground set $U\cup V$, where the hyperedges are the sets of the form $X\cup Y$, where $X\subseteq U$, $Y\subseteq V$, and $d(Y,c(X))=1$. This operation generalises natural matroid operations such as direct sums, $2$-sums, and proper amalgams.
@@
We are particularly interested in classes of gain-graphic matroids. These classes are analogous to classes of matroids representable by vectors. In a representable matroid the elements of the matroid are associated with vectors and in the gain-graphic case they are associated with elements from a group. In either case, we are essentially providing algebraic coordinates that specify the relations between matroid elements. Despite the symmetry of their roles in structural matroid theory, gain-graphic classes have not received as much attention as representable classes. The following conjectures appear in [FMN21].
Conjecture 1.1.
conj:finite-definable
@@ 2. Preliminaries
sec:prelims
We write $\mathbb{N}$ for the set of positive integers. If $n$ is in $\mathbb{N}$, we write $[n]$ for $\{1,2,\ldots,n\}$. We write $2^{U}$ for the power set of the set $U$. If $I$ is a set and $i$ is an element then we write $I+i$ for $I\cup\{i\}$. We regard each function as a set of ordered pairs. So if $\sigma\colon X\to Y$ is a function and $x$ is an element not in $X$ then $\sigma+(x,y)$ is the function with domain $X+x$ which takes each element of $x$ to its image under $\sigma$ and which takes $x$ to $y$. Graphs may have loops and parallel edges. If $G$ is a graph and $X$ is a set of edges, then $G[X]$ is the subgraph with $X$ as its set of edges. The vertices of $G[X]$ are exactly the vertices of $G$ that are incident with at least one edge in $X$. We will very frequently blur the distinction between sets of edges and subgraphs. For example, a cycle may be a set of edges or it may be a subgraph, according to which is more convenient for us.
@@ 2.1. Hypergraphs
The foundations of matroid theory can be found in Oxley [Oxl11]. A matroid is a hypergraph where the collection of hyperedges is non-empty and is closed downwards under subset containment, and furthermore, whenever $I$ and $J$ are hyperedges satisfying $|I|<|J|$, then there is an element $e\in J-I$ such that $I+e$ is a hyperedge. The hyperedges of a matroid are called independent sets. The dependent subsets are the subsets of the ground set that are not independent. A dependent subset that does not properly contain a dependent subset is a circuit. A (matroidal) loop is an element $e$ such that $\{e\}$ is a circuit. A coloop is an element that is in no circuit. A matroid is simple if every subset of size at most two is independent.
@@
2.2. Monadic second-order logic
@@ 4. $\bullet$
The collection of $\operatorname{CMSO}_{1}$-formulas that we construct without using any atomic formula of the form $|\cdot|\equiv p\bmod{q}$ is monadic second-order logic for hypergraphs, which we denote by $\operatorname{MSO}_{1}$. Let $\delta$ be a positive integer. We say that a $\operatorname{CMSO}_{1}$-formula is $\delta$-confined if it can be constructed without using a predicate of the form $|\cdot|\equiv p\bmod{q}$ where $q>\delta$. Note that a formula is $1$-confined if and only if it is an $\operatorname{MSO}_{1}$-formula.
@@ 2.3. Gain-graphic matroids
The rank of $F(\Omega)$ is obtained by subtracting the number of connected components in $\Omega$ that contain no unbalanced cycles from the number of vertices in $\Omega$. A line of a matroid is a rank-$2$ flat. The intersection of two distinct lines has rank at most one. A line is long if it contains at least four rank-$1$ flats.
@@
We can let $\{a,b,c,e\}$ and $\{x,y,z,e\}$ be sets such that any $3$-element subset of either is a circuit. We can also assume that no rank-$2$ flat contains both sets. Assume that $\{a,b,e\}$ is the edge-set of a balanced cycle in $\Omega$. Then $a$ and $b$ form a path of two edges, and $e$ must form a circuit with these edges. This is only possible if $\{a,b,c\}$ is also the set of edges in a balanced cycle, which means $c$ and $e$ are parallel edges. Therefore $\{a,b,c,e\}$ is the edge-set of a theta subgraph, but the cycle comprising $c$ and $e$ is not balanced, since $\{a,c,e\}$ is a circuit. Therefore this theta subgraph contains exactly two balanced cycles, and we have a contradiction. Exactly the same argument shows that no three edges from $\{a,b,c,e\}$ or $\{x,y,z,e\}$ form a balanced cycle. This means that $\{a,b,c\}$ is a bicycle that contains no balanced cycle. Subsequently $G[\{a,b,c\}]$ contains exactly two vertices. Let these vertices be $p$ and $q$. Then $e$ is not incident with any vertex not in $\{p,q\}$. The same argument shows that $G[\{x,y,z\}]$ has exactly two vertices, $s$ and $t$, and $e$ is not incident with any vertex not in $\{s,t\}$. If $\{s,t\}=\{p,q\}$ then there is a rank-$2$ flat that contains $\{a,b,c,x,y,z,e\}$, contrary to assumption. Therefore $\{s,t\}$ and $\{p,q\}$ have at most one vertex in common, which means that $e$ is incident with at most one vertex. Thus $e$ is a loop, as we claimed. ∎
@@ 2.4. Ultrapowers
Call these elements $h_{1},\ldots,h_{K}$. Suppose they generate a finite subgroup of $\Gamma^{\mathcal{U}}$, and let the size of this subgroup be $R$. This finiteness is witnessed by a finite sequence of equalities $\{s_{j}(h_{1},h_{1}^{-1},\ldots,h_{K},h_{K}^{-1})=\mathrm{Id}:j\in J\}$ (e.g. let these equalities imply all the equalities in the Cayley table of this group, thus forcing that any set satisfying those equalities is a quotient of this finite group). For each $j\in J$, let $S_{j}=\{i\in\mathbb{N}:s_{j}(g_{1,i},g_{1,i}^{-1},\ldots,g_{K,i},g_{K,i}^{-1})=\mathrm{Id}\}$. Then by the definition of $\sim_{\mathcal{U}}$, each $S_{j}$ is in $\mathcal{U}$, and by property (ii) of ultrafilters, $S=\bigcap_{j}S_{j}$ is also in $\mathcal{U}$, and is therefore nonempty. As in Remark 2.4, since the ultrafilter is non-principal, $S$ must be infinite. Fix an index $i$ in $S$ larger than $R$. Then $\{g_{1,i},\ldots,g_{K,i}\}$ satisfy all the $s_{j}$, so they must generate a subgroup of size at most $R$. But by definition of the sequence $g_{i,j}$, $\{g_{1,i},\ldots,g_{K,i}\}$ generates a subgroup of size at least $i$ for each $i$. Contradiction. The result follows. ∎
@@ 3. A Myhill-Nerode analogue
The models that we have in mind for coloured sums are versions of matroid sums. Standard matroid operations, such as $1$-, 22-, and $3$-sums can all be expressed as $C$-coloured sums, as can amalgams over finite sets.
To illustrate, we let $M$ and $N$ be matroids on the ground sets $U\cup\{p\}$ and $V\cup\{p\}$ respectively, where $U$, $V$, and $\{p\}$ are pairwise disjoint, and $p$ is not a loop or coloop in either matroid. We define the function $c\colon 2^{U}\to[3]$ as follows. If $X\subseteq U$ is dependent in $M$, then set $c(X)=1$. If $X$ is independent and $p$ is contained in the closure of $X$, then set $c(X)=2$. Otherwise, $X$ is independent and $p$ is not in the closure of $X$. In this case, set $c(X)$ to be $3$. Now $(U,c)$ is a $[3]$-coloured system. Let $d\colon 2^{V}\times[3]\to\{0,1\}$ be the function such that $(Y,i)$ is taken to $1$ if $Y$ is an independent subset of $N$ and either $i=3$, or $i=2$ and $Y$ does not span $p$. All other pairs are taken to 0. It is not difficult to check that the $[3]$-coloured sum $(U,c)\boxplus(V,d)$ is the hypergraph of independent sets of the $2$-sum of $M$ and $N$ along the basepoint $p$.
@@
For any prime number $p$, let $(U_{p},\mathcal{I}_{p})$ be a matroid isomorphic to the finite projective plane $\mathrm{PG}(2,p)$. Set $c_{p}$ to be the function which takes every dependent subset of $U_{p}$ to $1$ and every independent subset to $2$. Then $M_{p}=(U_{p},c_{p})$ is a $[2]$-coloured system. Because there are infinitely many prime numbers [Euclid]*Proposition 20, Lemma 3.7 tells us that there are distinct primes, $p_{1}$ and $p_{2}$, such that $M_{p_{1}}\sim_{\mathcal{M},[2]}M_{p_{2}}$ holds.
Let $N$ be a matroid $(V,\mathcal{I})$ that is isomorphic to $\mathrm{PG}(2,p_{1})$. Let $d\colon 2^{V}\times[2]\to\{0,1\}$ be the function taking $(Y,i)$ to $1$ when $Y$ is independent in $N$ and $i=2$. Other pairs are taken to 0. Then $M_{p_{i}}\boxplus(V,d)$ is isomorphic to the matroidal direct sum $\mathrm{PG}(2,p_{i})\oplus N$. Therefore $M_{p_{1}}\boxplus(V,d)$ is representable over $\mathrm{GF}(p_{1})$ [Oxl11]*Proposition 4.2.11. On the other hand, both $\mathrm{PG}(2,p_{1})$ and $\mathrm{PG}(2,p_{2})$ are isomorphic to minors of $M_{p_{2}}\boxplus(V,d)$ [Oxl11]*4.2.19. It follows from [Oxl11]*Proposition 3.2.4 and [Aig97]*Proposition 7.3 that if $M_{p_{2}}\boxplus(V,d)$ is representable over a field, then that field must simultaneously have subfields isomorphic to $\mathrm{GF}(p_{1})$ and $\mathrm{GF}(p_{2})$, an impossibility. To summarise, $M_{p_{1}}\boxplus(V,d)$ is representable and $M_{p_{2}}\boxplus(V,d)$ is not. Therefore $M_{p_{1}}\nsim_{\mathcal{M},[2]}M_{p_{2}}$, which is a contradiction. ∎
@@ 3. $\bullet$
prop:compatibility
Let $C$ be a finite set and let $\delta$ be a positive integer. Let $M=(U,c)$ be a $C$-coloured system. Let $\varphi$ be a $\delta$-confined $\operatorname{CMSO}_{1}$-formula. Set $S$ to be $\mathrm{Var}(\varphi)$ and let $\sigma$ be a function from $\mathrm{Free}(\varphi)$ to $2^{U}$. Let $\Pi=(V,d)$ be a $C$-coloured complement where we assume $U$ and $V$ are disjoint, and let $\tau$ be a function from $\mathrm{Free}(\varphi)$ to $2^{V}$. Set $E$ to be $U\cup V$ and define $\theta\colon\mathrm{Free}(\varphi)\to 2^{E}$ so that $\theta(Z_{s})=\sigma(Z_{s})\cup\tau(Z_{s})$ for every free variable $Z_{s}$. Let $N$ be the hypergraph $M\boxplus\Pi$. Then $\varphi$ is satisfied by $(N,\theta)$ if and only if $R_{S,C,\delta}(M,\varphi,\sigma)$ and $(\Pi,\tau)$ are sympathetic.

### v3

@@ 1. Introduction
In this article, we present a version of the Myhill-Nerode theorem for hypergraphs, which we will use to prove that certain classes of hypergraphs are not monadically definable. In particular, we will prove non-definability results when the hypergraphs are gain-graphic matroids. In the classical Myhill-Nerode theorem, $\mathcal{L}$ is a language of strings and two strings, $\mathbf{w}_{1}$ and $\mathbf{w}_{1}$, are equivalent relative to $\mathcal{L}$, if, for every possible string $\mathbf{z}$, either both of the concatenations $\mathbf{w}_{1}\mathbf{z}$ and $\mathbf{w}_{2}\mathbf{z}$ are in $\mathcal{L}$, or neither is. The language is regular if and only if this equivalence relation has finite index. To present a theorem of this type for hypergraphs, we need an operation for hypergraphs that plays the same role as concatenation in the classical theorem. Let $C$, $U$, and $V$ be finite sets where $U\cap V=\emptyset$. Let $c$ be a function from $2^{U}$ to $C$ and let $d$ be a function from $2^{V}\times C$ to $\{0,1\}$. We can glue $(U,c)$ and $(V,d)$ together to produce a hypergraph $(U,c)\boxplus(V,d)$ with the ground set $U\cup V$, where the hyperedges are the sets of the form $X\cup Y$ satisfying $X\subseteq U$, $Y\subseteq V$, and $d(Y,c(X))=1$. This operation generalises natural matroid operations such as direct sums, 22-sums, and proper amalgams.
@@
Courcelle’s Theorem for graphs [Cou90] provided much of the original motivation for studying the monadic second-order logic of graphs. There are matroidal analogues of this theorem which provide a similar motivation for wanting to know when a property of matroids can be defined in $\operatorname{CMSO}_{1}$ [FMN22, Hli03a, Kra12, Str11]. In particular, we are motivated to understand which minor-closed classes of matroids are definable in $\operatorname{CMSO}_{1}$. (This question is not interesting in the context of graphs, because the Robertson-Seymour Theorem shows that any minor-closed class of graphs has finitely many excluded minors and is therefore monadically definable.) For every group $\Gamma$ there is a corresponding class of $\Gamma$-gain-graphic matroids, just as for every field $\mathbb{F}$, there is a class of $\mathbb{F}$-representable matroids. In a representable matroid the elements of the matroid are associated with vectors over $\mathbb{F}$, and in the gain-graphic case they are associated with elements from $\Gamma$. In either case, we are essentially providing algebraic coordinates that specify the relations between matroid elements. The two types of classes play central and symmetric roles in structural matroid theory [GGW13, GNW24, KK82]. Despite this, gain-graphic classes have not received as much attention as representable classes. A result by Mayhew, Newman, and Whittle shows that the class of $\mathbb{F}$-representable matroids is $\operatorname{CMSO}_{1}$-definable if and only if $\mathbb{F}$ is finite. The following conjectures, from [FMN21], were made by analogy with this result.
Conjecture $1.$
@@ 2. Preliminaries
We write $\mathbb{N}$ for the set of positive integers. If $n$ is in $\mathbb{N}$, we write $[n]$ for $\{1,2,\ldots,n\}$. We write $2^{U}$ for the power set of the set $U$. If II is a set and $i$ is an element then we write $I+i$ for $I\cup\{i\}$. We regard each function as a set of ordered pairs. So if $\sigma\colon X\to Y$ is a function and $x$ is an element not in $X$ then $\sigma+(x,y)$ is the function with domain $X+x$ which takes each element of $x$ to its image under $\sigma$ and which takes $x$ to $y$. Graphs may have loops and parallel edges. If $G$ is a graph and $X$ is a set of edges, then $G[X]$ is the subgraph with $X$ as its set of edges. The vertices of $G[X]$ are exactly the vertices of $G$ that are incident with at least one edge in $X$. We will very frequently blur the distinction between sets of edges and subgraphs. For example, a cycle may be a set of edges or it may be a subgraph, according to which is more convenient for us.
@@ 2.1. Hypergraphs
The foundations of matroid theory can be found in Oxley [Oxl11]. A matroid is a hypergraph where the collection of hyperedges is non-empty and is closed downwards under subset containment, and furthermore, whenever II and $J$ are hyperedges satisfying $|I|<|J|$, then there is an element $e\in J-I$ such that $I+e$ is a hyperedge. The hyperedges of a matroid are called independent sets. The dependent subsets are the subsets of the ground set that are not independent. A dependent subset that does not properly contain a dependent subset is a circuit. A (matroidal) loop is an element $e$ such that $\{e\}$ is a circuit. A coloop is an element that is in no circuit. A matroid is simple if every subset of size at most two is independent.
@@
$2.$ Monadic second-order logic
@@ 4. $\bullet$
The collection of $\operatorname{CMSO}_{1}$-formulas that we construct without using any atomic formula of the form $|\cdot|\equiv p\bmod{q}$ is monadic second-order logic for hypergraphs, which we denote by $\operatorname{MSO}_{1}$. Let $\delta$ be a positive integer. We say that a $\operatorname{CMSO}_{1}$-formula is $\delta$-confined if it can be constructed without using a predicate of the form $|\cdot|\equiv p\bmod{q}$ where $q>\delta$. Note that a formula is 11-confined if and only if it is an $\operatorname{MSO}_{1}$-formula.
@@ 2.3. Gain-graphic matroids
The rank of $F(\Omega)$ is obtained by subtracting the number of connected components in $\Omega$ that contain no unbalanced cycles from the number of vertices in $\Omega$. A line of a matroid is a rank-22 flat. The intersection of two distinct lines has rank at most one. A line is long if it contains at least four rank-11 flats.
@@
We can let $\{a,b,c,e\}$ and $\{x,y,z,e\}$ be sets such that any 33-element subset of either is a circuit. We can also assume that no rank-22 flat contains both sets. Assume that $\{a,b,e\}$ is the edge-set of a balanced cycle in $\Omega$. Then $a$ and $b$ form a path of two edges, and $e$ must form a circuit with these edges. This is only possible if $\{a,b,c\}$ is also the set of edges in a balanced cycle, which means $c$ and $e$ are parallel edges. Therefore $\{a,b,c,e\}$ is the edge-set of a theta subgraph, but the cycle comprising $c$ and $e$ is not balanced, since $\{a,c,e\}$ is a circuit. Therefore this theta subgraph contains exactly two balanced cycles, and we have a contradiction. Exactly the same argument shows that no three edges from $\{a,b,c,e\}$ or $\{x,y,z,e\}$ form a balanced cycle. This means that $\{a,b,c\}$ is a bicycle that contains no balanced cycle. Subsequently $G[\{a,b,c\}]$ contains exactly two vertices. Let these vertices be $p$ and $q$. Then $e$ is not incident with any vertex not in $\{p,q\}$. The same argument shows that $G[\{x,y,z\}]$ has exactly two vertices, $s$ and $t$, and $e$ is not incident with any vertex not in $\{s,t\}$. If $\{s,t\}=\{p,q\}$ then there is a rank-22 flat that contains $\{a,b,c,x,y,z,e\}$, contrary to assumption. Therefore $\{s,t\}$ and $\{p,q\}$ have at most one vertex in common, which means that $e$ is incident with at most one vertex. Thus $e$ is a loop, as we claimed. ∎
@@ 2.4. Ultrapowers
Call these elements $h_{1},\ldots,h_{K}$. Suppose they generate a finite subgroup of $\Gamma^{\mathcal{U}}$, and let the size of this subgroup be $R$. This finiteness is witnessed by a finite sequence of equalities $\{s_{j}(h_{1},h_{1}^{-1},\ldots,h_{K},h_{K}^{-1})=\mathrm{Id}:j\in J\}$ (e.g. let these equalities imply all the equalities in the Cayley table of this group, thus forcing that any set satisfying those equalities is a quotient of this finite group). For each $j\in J$, let $S_{j}=\{i\in\mathbb{N}:s_{j}(g_{1,i},g_{1,i}^{-1},\ldots,g_{K,i},g_{K,i}^{-1})=\mathrm{Id}\}$. Then by the definition of $\sim_{\mathcal{U}}$, each $S_{j}$ is in $\mathcal{U}$, and by property (ii) of ultrafilters, $S=\bigcap_{j}S_{j}$ is also in $\mathcal{U}$, and is therefore nonempty. As in Remark 2.4, since the ultrafilter is non-principal, $S$ must be infinite. Fix an index $i$ in $S$ larger than $R$. Then $\{g_{1,i},\ldots,g_{K,i}\}$ satisfy all the $s_{j}$, so they must generate a subgroup of size at most $R$. But by definition of the sequence $g_{i,j}$, $\{g_{1,i},\ldots,g_{K,i}\}$ generates a subgroup of size at least $i$ for each ii. Contradiction. The result follows. ∎
@@ 3. A Myhill-Nerode analogue
The models that we have in mind for coloured sums are versions of matroid sums. Standard matroid operations, such as 11-, 22-, and 33-sums can all be expressed as $C$-coloured sums, as can amalgams over finite sets.
To illustrate, we let $M$ and NN be matroids on the ground sets $U\cup\{p\}$ and $V\cup\{p\}$ respectively, where $U$, $V$, and $\{p\}$ are pairwise disjoint, and $p$ is not a loop or coloop in either matroid. We define the function $c\colon 2^{U}\to[3]$ as follows. If $X\subseteq U$ is dependent in $M$, then set $c(X)=1$. If $X$ is independent and $p$ is contained in the closure of $X$, then set $c(X)=2$. Otherwise, $X$ is independent and $p$ is not in the closure of $X$. In this case, set $c(X)$ to be 33. Now $(U,c)$ is a $[3]$-coloured system. Let $d\colon 2^{V}\times[3]\to\{0,1\}$ be the function such that $(Y,i)$ is taken to 11 if $Y$ is an independent subset of NN and either $i=3$, or $i=2$ and $Y$ does not span pp. All other pairs are taken to 0. It is not difficult to check that the $[3]$-coloured sum $(U,c)\boxplus(V,d)$ is the hypergraph of independent sets of the 22-sum of $M$ and NN along the basepoint pp.
@@
For any prime number $p$, let $(U_{p},\mathcal{I}_{p})$ be a matroid isomorphic to the finite projective plane $\mathrm{PG}(2,p)$. Set $c_{p}$ to be the function which takes every dependent subset of $U_{p}$ to 11 and every independent subset to 22. Then $M_{p}=(U_{p},c_{p})$ is a $[2]$-coloured system. Because there are infinitely many prime numbers [Euclid]*Proposition 20, Lemma 3.7 tells us that there are distinct primes, $p_{1}$ and $p_{2}$, such that $M_{p_{1}}\sim_{\mathcal{M},[2]}M_{p_{2}}$ holds.
Let NN be a matroid $(V,\mathcal{I})$ that is isomorphic to $\mathrm{PG}(2,p_{1})$. Let $d\colon 2^{V}\times[2]\to\{0,1\}$ be the function taking $(Y,i)$ to 11 when $Y$ is independent in NN and $i=2$. Other pairs are taken to 0. Then $M_{p_{i}}\boxplus(V,d)$ is isomorphic to the matroidal direct sum $\mathrm{PG}(2,p_{i})\oplus N$. Therefore $M_{p_{1}}\boxplus(V,d)$ is representable over $\mathrm{GF}(p_{1})$ [Oxl11]*Proposition 4.2.11. On the other hand, both $\mathrm{PG}(2,p_{1})$ and $\mathrm{PG}(2,p_{2})$ are isomorphic to minors of $M_{p_{2}}\boxplus(V,d)$ [Oxl11]*4.2.19. It follows from [Oxl11]*Proposition 3.2.4 and [Aig97]*Proposition 7.3 that if $M_{p_{2}}\boxplus(V,d)$ is representable over a field, then that field must simultaneously have subfields isomorphic to $\mathrm{GF}(p_{1})$ and $\mathrm{GF}(p_{2})$, an impossibility. To summarise, $M_{p_{1}}\boxplus(V,d)$ is representable and $M_{p_{2}}\boxplus(V,d)$ is not. Therefore $M_{p_{1}}\nsim_{\mathcal{M},[2]}M_{p_{2}}$, which is a contradiction. ∎
@@ 3. $\bullet$
Let $C$ be a finite set and let $\delta$ be a positive integer. Let $M=(U,c)$ be a $C$-coloured system. Let $\varphi$ be a $\delta$-confined $\operatorname{CMSO}_{1}$-formula. Set $S$ to be $\mathrm{Var}(\varphi)$ and let $\sigma$ be a function from $\mathrm{Free}(\varphi)$ to $2^{U}$. Let $\Pi=(V,d)$ be a $C$-coloured complement where we assume $U$ and $V$ are disjoint, and let $\tau$ be a function from $\mathrm{Free}(\varphi)$ to $2^{V}$. Set $E$ to be $U\cup V$ and define $\theta\colon\mathrm{Free}(\varphi)\to 2^{E}$ so that $\theta(Z_{s})=\sigma(Z_{s})\cup\tau(Z_{s})$ for every free variable $Z_{s}$. Let NN be the hypergraph $M\boxplus\Pi$. Then $\varphi$ is satisfied by $(N,\theta)$ if and only if $R_{S,C,\delta}(M,\varphi,\sigma)$ and $(\Pi,\tau)$ are sympathetic.